# Estimates for trilinear and quadrilinear character sums

### v2

@@ 1.1. Set-up
Motivated by various applications to analytic number theory, estimates for character sums received a lot of attentions in the past decades. In many situations, the variables are supported over some additively structured sets, such as intervals. But the difficulty can vary significantly since the weights might be arbitrary. This paper aims to study two kinds of multilinear character sums with arbitrary weights.
Throughout this paper, denote by $p$ a large odd prime, and by $\chi$ a non-trivial multiplicative character modulo $p$. Take two integers $a,b$ with $p\nmid ab$. For $K,L,M,N\geqslant 1$, we consider the trilinear character sum
@@
where $\bm{\alpha}=(\alpha_{m})$ or $(\alpha_{\ell,m})$ and $\bm{\beta}=(\beta_{k,n})$. Note that one obtains (1.1) if taking $L=1$ and $\alpha_{1,m}=\alpha_{m}$ in (1.2). No confusion on the definitions of $\bm{\alpha}$ arises since one is for the trilinear sum, and the other is for the quadrilinear one.
@@ 1.2. Related bilinear sums
The above studies on character sums over sumsets can date back to Vinogradov [Vi54, Chapter V, Exercise 8.c] on the following bilinear form
@@
Although this bound is widely known, its full derivation is not easy to find in the literature, however it can be found as a short proof of [SS17, Equation (1.4)]. Despite a very short and elementary proof, the bound (1.4) has never been improved in full generality. On the other hand, Karatsuba [Ka92] (see also [Ka93, Chapter VIII, Problem 9]) can prove, that if
@@
holds with some $\kappa>0$, depending only on $\eta$. The range (1.5) reveals the Pólya–Vinogradov threshold even when summing over arbitrary subsets. Similar phenomenon can also be found in [Xi23]. We note that the proof of (1.6), in the range (1.5), does not seem to be present in literature, and a concise proof with an explicit upper bound, is provided in Appendix A to convince cautious readers.
It is worthwhile to point out that $\mathfrak{B}(\bm{\alpha},\bm{\beta})$ has received considerable attentions in recent years due to its connection with the Paley graph conjecture. It is conjectured, for instance with just constant weights $\alpha_{m}=\beta_{n}=1$, that
@@ 1.3. Related trilinear and quadrilinear sums
with bounded weights $(\alpha_{r})$, $(\beta_{u})$ and $(\gamma_{r,s,u})$. It is shown in [FS23], that for any fixed $\eta>0$ and $x\geqslant p^{\frac{1}{8}+\eta}$, we have
@@ 2.1. Bounds of multilinear character sums
$|\mathfrak{Q}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|_{2}(LM)(KN)^{\frac{1}{2}}p^{o(1)}\cdot\Delta_{3},$
@@
$|\mathfrak{Q}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|(LM)(KN)^{\frac{1}{2}}p^{-\kappa}.$
@@ 2.2. Consequences and applications
To illustrate Theorem 2.11, choose $M=N\sim p^{\frac{1}{5}+\delta}$ (with $\delta>0$ very small) and ${\mathcal{M}}$ and ${\mathcal{N}}$ subsets of $[N,2N]$ such that #,#$\#{\mathcal{M}},\#{\mathcal{N}}\geqslant\delta N$. Then, for sufficiently large $p$, there exists a positive integer $k\leqslant p^{\frac{1}{5}}$ and $(m,n)\in{\mathcal{M}}\times{\mathcal{N}}$ such that $mn+k$ is a quadratic residue modulo $p$. Recall that there must exist a square of an integer between $t$ and $t+O(\sqrt{t})$ for all $t>1$, and that essentially, nothing better than the upper bound $p^{\frac{1}{4}+o(1)}$ is proved for the distance between two consecutive quadratic residues modulo $p$. The above illustration shows that one can do much better if squares are replaced by quadratic residues modulo $p$.
@@
$\alpha_{\ell,m}\beta_{k,n}\chi(\ell\overline{m}-\overline{k}n)=\left(\alpha_{\ell,m}\overline{\chi}(m)\right)\left(\beta_{k,n}\overline{\chi}(k)\right)\chi(k\ell-mn),$
@@ 3. Preliminaries
3.1. Notation and conventions
We also use ${\mathbb{N}}$ to denote that set of positive integers.
As usual,
(3.2) $\tau(k)=\#\{d\in{\mathbb{N}}:\leavevmode\nobreak\ d\mid k\}$
$\widehat{f}(y)=\int_{\mathbb{R}}f(x)\mathrm{e}(-yx)\mathrm{d}x$
with $\mathrm{e}(z)=\exp(2\pi iz)$.
@@ 3.3. Bounds of some GCD sums
Denote by $G$ the above quantity in question. Writing $g_{1}=(a_{1},b_{1})$ and $g_{2}=(a_{2},b_{2})$ we have the inequality

### v3

@@ 1.1. Set-up
Motivated by various applications to analytic number theory, estimates for character sums received a lot of attention in the past decades. In many situations, the variables are supported over some additively structured sets, such as sets of consecutive integers. But the difficulty can vary significantly since the weights might be arbitrary. This paper aims to study two kinds of multilinear character sums with arbitrary weights.
Throughout this paper, denote by $p$ a large prime, and by $\chi$ a non-trivial multiplicative character modulo $p$. Take two integers $a,b$ with $p\nmid ab$. For $K,L,M,N\geqslant 1$, we consider the trilinear character sum
@@
where $\bm{\alpha}=(\alpha_{m})$ or $(\alpha_{\ell,m})$ and $\bm{\beta}=(\beta_{k,n})$ are some complex weights. Note that one obtains (1.1) if taking $L=1$ and $\alpha_{1,m}=\alpha_{m}$ in (1.2). No confusion on the definitions of $\bm{\alpha}$ arises since one is for the trilinear sum, and the other is for the quadrilinear one.
@@ 1.2. Related bilinear sums
The above studies on character sums over sumsets can be dated back to Vinogradov [Vi54, Chapter V, Exercise 8.c] on the following bilinear form
@@
Although this bound is widely known, its full derivation is not easy to find in the literature, however it can be found as a short proof of [SS17, Equation (1.4)]. Despite a very short and elementary proof, the bound (1.4) has never been improved in full generality. However, Karatsuba [Ka92] (see also [Ka93, Chapter VIII, Problem 9]) has proved, that if
@@
holds with some $\kappa>0$, depending only on $\eta$. The range (1.5) reveals the Pólya–Vinogradov threshold even when summing over arbitrary subsets. A similar phenomenon can also be found in [Xi23]. We note that the proof of (1.6), in the range (1.5), does not seem to be present in literature, and a concise proof with an explicit upper bound, is provided in Appendix A to convince cautious readers.
It is worthwhile to point out that $\mathfrak{B}(\bm{\alpha},\bm{\beta})$ has received considerable attention in recent years due to its connection with the Paley graph conjecture. It is conjectured, for instance with just constant weights $\alpha_{m}=\beta_{n}=1$, that
@@ 1.3. Related trilinear and quadrilinear sums
with bounded complex weights $\bm{\alpha}=(\alpha_{r})$, $\bm{\beta}=(\beta_{u})$ and $\bm{\gamma}=(\gamma_{r,s,u})$. It is shown in [FS23], that for any fixed $\eta>0$ and $x\geqslant p^{\frac{1}{8}+\eta}$, we have
@@
1.4. Notation and conventions
We also use ${\mathbb{N}}$ to denote that set of positive integers.
As usual,
(1.8) $\tau(k)=\#\{d\in{\mathbb{N}}:\leavevmode\nobreak\ d\mid k\}$
$\widehat{f}(y)=\int_{\mathbb{R}}f(x)\mathrm{e}(-yx)\mathrm{d}x$
with $\mathrm{e}(z)=\exp(2\pi iz)$.
@@ 2.1. Bounds of multilinear character sums
$|\mathfrak{Q}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|_{2}LM(KN)^{\frac{1}{2}}p^{o(1)}\cdot\Delta_{3},$
@@
$|\mathfrak{Q}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|LM(KN)^{\frac{1}{2}}p^{-\kappa}.$
@@ 2.2. Consequences and applications
To illustrate Theorem 2.11, choose $M=N\sim p^{\frac{1}{5}+\delta}$ (with $\delta>0$ very small) and ${\mathcal{M}}$ and ${\mathcal{N}}$ subsets of $[N,2N]$ such that #,#$\#{\mathcal{M}},\#{\mathcal{N}}\geqslant\delta N$. Then, for sufficiently large $p$, there exists a positive integer $k\leqslant p^{\frac{1}{5}}$ and $(m,n)\in{\mathcal{M}}\times{\mathcal{N}}$ such that $mn+k$ is a quadratic residue modulo $p$. Recall that there must exist a square of an integer between $t$ and $t+O(\sqrt{t})$ for all $t>1$ (namely, $\lceil\sqrt{t}\rceil^{2}$ is such a square) and that essentially, nothing better than the upper bound $p^{\frac{1}{4}+o(1)}$ is proved for the distance between two consecutive quadratic residues modulo $p$. The above illustration shows that one can do much better if squares are replaced by quadratic residues modulo $p$.
@@
$\alpha_{\ell,m}\beta_{k,n}\chi(\ell\overline{m}-\overline{k}n)=\left(\alpha_{\ell,m}\overline{\chi}(m)\right)\left(\beta_{k,n}\overline{\chi}(k)\right)\chi(k\ell-mn)$
@@ 3.2. Bounds of some GCD sums
Denote by $G$ the above quantity in question. Writing $g_{1}=\gcd(a_{1},b_{1})$ and $g_{2}=\gcd(a_{2},b_{2})$ we have the inequality